A bakery offers 32 types of cakes and cookies. If there are 6 more types of cookies than cakes, how many types of cakes are available?
Answer by ptaylor(2198)

Let x=number of types of cakes
Then x+6=number of types of cookies

Now we are told that:
x+x+6=32 subtract 6 from both sides
x+x+6-6=32-6 collect like terms
2x=26 divide both sides by 2
x=13---------------------number of types of cakes
x+6=13+6=19------------------number of types of cookies
CK
13+19=32
32=32
also 19 is 6 more than 13

Another way:
Let x=number of types of cakes
And y=number of types of cookies
Now we are told that x+y=32--------------------------eq1
We are also told that y=x+6--------------------------eq2
Rearrange eq2 by subtracting x from both sides and we get:
y-x=6--------------------------------eq2
Add eq1 and eq2:
2y=38 divide both sides by 2
y=19--------------------number of types of cookies
substitute y=19 into eq1
x+19=32 subtract 19 from both sides
x+19-19=32-19 collect like terms
x=13---------------------------------number of types of cakes
Hope this helps---ptaylor